OpenStudy (anonymous):

Hey, can someone explain an equation to me step by step? Ill fan and give a Medal. I have to solve this using substitution, and I'm having a bit of difficulty. -4 + y = 6 -5x -y = 21 I know the answer is ( -3, -6) Im just having difficulty understanding how you get that answer

OpenStudy (anonymous):

@Jlockwo3

OpenStudy (anonymous):

Do I just change both of the equations to solve for y?

OpenStudy (anonymous):

pick the one that does not have both x and y in it

OpenStudy (anonymous):

y= 6+4

OpenStudy (anonymous):

I have to solve them using substitution, not graphing

OpenStudy (anonymous):

then do what we did last time, where we plugged that y= into the equation that has both x and y in it

OpenStudy (anonymous):

\[-5x - (10)= 21\]

OpenStudy (anonymous):

my bad! she requested by presence

OpenStudy (anonymous):

So.. y = 6 + 4x -5x - (6 + 4x) =21

OpenStudy (anonymous):

just want to clarify...did you mis type above? was it -4 +y =6 ?

OpenStudy (anonymous):

if that is correct. y = 6+4 then -5x - (10) =21

OpenStudy (anonymous):

Oh gosh, Im so sorry. Yes it was. Its 4x + y =6

OpenStudy (anonymous):

ok. that changes things... but its not as bad as you think.

OpenStudy (anonymous):

change that to y = 6 + 4x and then plug that in, and you were correct in your assumption as noted above

OpenStudy (anonymous):

So would the equation then be 6 + 9x = 21

OpenStudy (anonymous):

\[-4x + y = 6 --> y= 6+4x---> -5x- (6+4x)= 21--> -5x - 6 -4x=21\]

OpenStudy (anonymous):

you have to distribute that neg sign...

OpenStudy (anonymous):

\[-5x-(6+4x)=21 --> -5x -6 -4x=21\]

OpenStudy (anonymous):

Oh, I get it now. Thanks hun!!!

OpenStudy (anonymous):

\[-9x = 21+6\]

OpenStudy (anonymous):

np

OpenStudy (anonymous):

jeez he was rude to me! did he not see that you asked me to come? lol

OpenStudy (anonymous):

and he was wrong too lol

OpenStudy (anonymous):

TO be honest, my computer has been acting so stupid that the only thing I saw was one second he had made like 10 comments trying to get me to solve this through graphing instead of substitution then next thing I know, all his comment disappeared

OpenStudy (anonymous):

he got mad that i started helping you

OpenStudy (anonymous):

Goodness, really? He must be new then, usually the members on here try helping each other out explaining the problem at hand instead of fighting over it. He must've really wanted a medal.

OpenStudy (anonymous):

lol